3d^2y/dx^2+dy/dx-14y=0 ; y=(0) ,y'=(0)=-1


1
Expert's answer
2022-04-12T16:12:42-0400

"3y''+y'-14y=0,"

"y=c_1e^{-\frac{7x}3}+c_2e^{2x},"

"y(0)=0: c_1=-c_2,"

"y'(0)=-1: c_2=\frac 76 c_1-\frac 12,"

"c_1=\frac3{13},~c_2=-\frac 3{13},"

"y=\frac 3{13}e^{-\frac{7x}3}-\frac 3{13}e^{2x}."